Solve the following algebraically. You must show all your work.
Learn how to type math roots and fractions by clicking on link in the assignment list. Alternately, you may type as cuberoot(x) and show raising to the nth power as ^n, like x 3 is typed x^3.
1/4=3- 2x-1/x+2

Answer by stanbon(48568)

(1/4) =3 - [(2x-1)/(x+2)]
[(2x-1)/(x+2)] = 3 - (1/4)
[(2x-1)/(x+2)] = 11/4
Cross-multiply to get:
8x-4 = 11x+22
3x = -26
x = -26/3
==============
Cheers,
Stan H.